Assuming that the oxygen gas is collected at P = 755 mm Hg and T = 305 K with the chemical equation 2KClO3(s) → 2KCl(s) + 3O2(g), how many liters of oxygen gas are produced when 294 g of KClO3 completely react as used in the ignition of fireworks? Solve to get the correct answer


1
Expert's answer
2021-11-08T08:23:49-0500

Solution:

Calculate the moles of KClO3:

The molar mass of KClO3 is 122.55 g/mol

Hence,

(294 g KClO3) × (1 mol KClO3 / 122.55 g KClO3) = 2.399 mol KClO3 = 2.4 mol KClO3


Balanced chemical equation:

2KClO→ 2KCl(s) + 3O2(g)

According to stoichiometry:

2 mol of KClO3 produces 3 mol of O2

Thus, 2.4 mol of KClO3 produces:

(2.4 mol KClO3) × (3 mol O2 / 2 mol KClO3) = 3.6 mol O2


The ideal gas law can be used: PV = nRT

To find the volume of O2, solve the equation for V:

V = nRT/P

V(O2) = (3.6 mol × 62.3637 L mmHg mol-1 K-1 × 305 K) / (755 mmHg) = 90.6958 L = 90.7 L

V(O2) = 90.7 L


Answer: 90.7 liters of oxygen gas are produced.